40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 5 18:57:10 2011, Jamie, Update, SUS, More normalization of all sus controllers 
    Reply  Tue Jul 5 21:26:51 2011, rana, Update, SUS, More normalization of all sus controllers Screenshot-1.png
       Reply  Wed Jul 6 02:12:36 2011, kiwamu, Update, SUS, Re : More normalization of all sus controllers 
          Reply  Wed Jul 6 10:35:35 2011, Jamie, Update, SUS, Re : More normalization of all sus controllers 
             Reply  Wed Jul 6 15:32:32 2011, Jamie, Update, SUS, Re : More normalization of all sus controllers 
       Reply  Wed Jul 6 11:45:20 2011, Jamie, Update, SUS, More normalization of all sus controllers 
Message ID: 4941     Entry time: Tue Jul 5 18:57:10 2011     Reply to this: 4942
Author: Jamie 
Type: Update 
Category: SUS 
Subject: More normalization of all sus controllers 

Based on Rana's comment I have gone through and moved all of the corner frequencies for the high pass filters in the SUS damping controllers to 30 Hz.  I did this for all optics (MC1, MC1, MC3, BS, ITMX, ITMY, PRM, SRM, ETMX, ETMY) all degrees of freedom (POS, PIT, YAW, SIDE).

Rana also suggested I turn off all of the BounceRoll filters until we get a chance to tune those individually for all the suspensions.

Finally, I normalized the MC SUSXXX filter banks to look just like all the other suspensions.

All damping filter banks for all degrees of freedom for all single suspensions should all be the same now (modulo the differences in the BounceRoll filters, which are now turned off).

ELOG V3.1.3-